# “There Has To Be a Lot That We’re Missing”: Moderating AI-Generated Content on Reddit

### v6

@@ 4.1.3. Governance Concerns
Participants also objected to AIGC because it made the job of moderating their communities more difficult. We heard from eight participants who saw AIGC being used for malicious purposes and were concerned that existing moderation techniques would struggle to handle it. These moderators spoke of pre-existing problems, such as spam and harassment, that are more difficult to manage when they are amplified in scale or sophistication by AIGC. While AIGC may be less common than other moderation challenges, it can still be quite disruptive:
@@
Five participants, who moderated the largest subreddits, spoke of AIGC being used in broad attacks on their communities. These attacks often involved automated accounts posting large amounts of spam. These attacks may aim to disrupt the functioning of a community through sheer volume, as we heard from one moderator who experienced a ChatGPT-powered spam attack that required the help of Reddit staff to resolve:
@@
Other times these attacks aim to deceive users as a part of influence operations, as we heard from a moderator of r/news, one of the platform’s largest subreddits, who shared their experience dealing with, “a highly sophisticated deployment, an attempt to sway a narrative by utilizing accounts that had had their karma bolstered by automated commenting.” This participant references Reddit karma666https://support.reddithelp.com/hc/en-us/articles/204511829-What-is-karma, a reputation points system that communities can use to limit who can post, noting that AIGC can be used to bypass this system through a practice known as karma-farming, in which automated accounts make frequent posts in order to rapidly accumulate karma. Karma-farming is not a new practice, but this moderator thought that AIGC made it harder to detect, and thus more effective. Beyond karma-farming, we also heard about AIGC being used in other forms of deception:
@@
Finally, we heard concerns about AI making it easier to post objectionable content. One participant who moderated r/Zelda shared that they had seen AI used to produce specific types of objectionable content that their community had already banned, such as commercial or NSFW content. They noted that while this type of content was not new, “it seemed to be new territory for people to expedite these things via AI.”
@@ 4.1.4. Potential Benefits of AIGC
Three moderators mentioned that AI could help non-native English-speakers improve their posts. We heard from a moderator of r/changemyview who thought that, “AI as a tool is going to be useful…it’s useful for people who are poor writers, useful for people who aren’t strong English speakers. And when they need help, AI can fill in those gaps,” though they noted that it was important that such a contribution was still, “sufficiently their words”. Of course, the definition of “sufficient” will be context dependent and may change over time as norms around AI use become more established.
One last benefit that we heard was that AIGC could increase engagement in an inactive conversation. For example, one moderator speculated about the potential benefits of AIGC for platforms and provided a hypothetical example:
@@ 4.2. Findings: Online Communities’ Responses to AIGC (RQ2)
To counter the perceived threats of AIGC, moderators have enacted specific changes in their communities. First, our participants emphasized collaborating with their community members to clarify norms about the use of AIGC. Once clarified, these norms could be codified into rules, which moderators were then tasked with enforcing.
@@ 4.2.1. Clarifying Community Norms
Almost all participants mentioned eventually codifying community norms into public subreddit rules. Some communities used rules that explicitly banned AIGC or bots, while others used more general rules (such as rules against plagiarism) as effective AIGC bans. Participants shared mixed results about the efficacy of these rules. We heard from seven participants that the volume of AIGC had decreased in response to new rules, including from a moderator of r/AskHistorians who said, “People are using [AIGC] a little bit less now that they know what the rules are. They know that people don’t necessarily want it.” However, a moderator of r/explainlikeimfive expressed skepticism about the efficacy of rules: “I think it may stop some of the average users (to the extent that they actually read the rules, which is a reddit-wide problem.) Bots obviously don’t read rules.” As this participant suggested, explicit rules can guide the behavior of good-faith users, but might not deter bad actors.
@@ 4.2.2. Enforcing Community Rules
Our participants used a variety of strategies to try to make enforcement of these rules less burdensome. Seven moderators mentioned sharing the work with community members, like a moderator of r/news who noted, “The first line of defense is the users themselves.” However, this runs the risk of causing incivility between members in the case of false accusations, as mentioned in Section 4.1.2.
Four moderators spoke of their reliance on platform tools, especially automoderator777https://www.reddit.com/wiki/automoderator/, a programmable bot that scans and performs actions on all new posts. Such automated tools lessened the manual effort required by moderators, but they are only as effective as the detection heuristics programmed into them, which, as we discuss in the next section, can be flawed. We heard from moderators that such tools are not enough to catch all AIGC, and in cases of extreme post volume, such as during coordinated bot attacks, moderators sometimes sought help from Reddit platform admins:
@@ 4.3. Findings: Methods for Detecting AIGC (RQ2)
Our participants shared both custom heuristics and technical solutions that they use to detect AIGC. Even if participants thought their detection techniques mostly work now, they worry that their tools will be less effective as generative AI improves:
@@ 4.3.2. Technical Tools
Though they acknowledged such tools’ unreliability, most participants used automated approaches to detect AIGC, including both third-party, black-box detection tools and community-developed solutions. Participants spoke of programming the detection heuristics mentioned above into tools such as automoderator, which apply the heuristics to every new post and flag anything suspect. Six participants mentioned manually copying posts into third-party detection tools to see if they were AIGC. A moderator of r/writingprompts reported that they had tested, but never deployed, a homespun bot that would apply third-party detectors automatically to all new posts.
@@ 5. Discussion
Our qualitative investigation offers two main research contributions: (1) a description of moderators’ attitudes towards AIGC in subreddits that restrict its use and (2) insights into how these moderators and their communities have responded to the challenges of AIGC. On the whole, our findings align with those of recent studies into the impact of generative AI on online communities (burtch_consequences_2024; wei2024understandingimpactaigenerated), which emphasize that the social value of online communities is more important than ever in the age of generative AI. Platform designers and community stewards should focus on ways to increase this social value, and one way to do this is by empowering communities to take their own stance on AIGC.
As generative AI usage grows, we place our work in dialog with Seering’s call to the CSCW research community to center community self-moderation (Seering20), especially their second “guiding question”: “What are the processes of context-sensitivity in online community moderation, and how might they be better supported?” We believe our findings demonstrate that allowing communities to make context-sensitive decisions about AIGC is central to their self-determination and should be an important topic for CSCW research. To offer implications for design, we expand below on the values that shape community attitudes towards AIGC as well the importance of clear and enforceable community norms.
@@ 5.1. Community Values Shape Attitudes Towards AIGC
Our study reveals a spectrum of moderator attitudes that vary with the purpose and size of communities. We expect certain objections to AIGC will lessen over time, while others will be more persistent.
@@ 5.1.1. Community Types and AIGC Attitudes
Our participants’ concerns about AIGC were largely based on the types of communities they moderated. Accordingly, it is useful to consider our results through the lens of several online community types from the Social Computing literature, such as knowledge communities (jenkins_introduction_2006), empathic communities (preece_empathic_1999), communities of practice (lave1991situated), and the five Community Archetypes presented by Prinster et al. (Prinster24): Topical Q&A, Learning & Perspective Broadening, Social Support, Content Generation, and Affiliation with an Entity. We suspect that AIGC’s threats to the social value of a community (Section 4.1.2) would be the most concerning to communities that emphasize authenticity (Smith23), like empathic communities (preece_empathic_1999), or communities that fit the Social Support archetype. Moderators of knowledge communities (jenkins_introduction_2006) like r/AskHistorians and r/explainlikeimfive, which also fit the Topical Q&A archetype, also shared this concern, though they additionally took issue with AIGC’s inaccuracy (Section 4.1.1). Arguably the clearest objections came from moderators of communities of practice (lave1991situated) like r/itookapicture and r/WritingPrompts, which fit the Content Generation archetype. These communities shared both content quality and social dynamics concerns about AIGC, though we expect new communities of practice will emerge specifically devoted to AI use888https://www.reddit.com/r/aiArt/. Moderators of larger, more impersonal communities that fit the Learning & Perspective Broadening archetype, such as r/news, r/CryptoCurrency, and r/politics, primarily voiced governance concerns (Section 4.1.3). We suspect this is because the size of these communities make them a target of bad actors. Finally, we primarily heard ideological objections from the communities in our sample that fit the Affiliation with an Entity archetype. We suspect this is because in our sample these are Content Generation communities (r/WoW, r/Zelda), and it remains to be seen if communities affiliated with non-media entities, like geographical places or organizations, would share similar concerns.
On the whole, our participants rejected AIGC because they felt the main purpose of their communities was to share posts created by humans—this was true across all of the community types that we encountered (of course, remember that our sampling strategy focused on communities with concerns, not representativeness). Future work should further explore the relationship between community type and attitude towards AIGC. For example, it would be interesting to see, quantitatively, whether different types of communities tend to create different types of rules about AIGC.
@@ 5.1.2. Attitudes May Shift Over Time
Our findings indicate that community norms about AIGC are still in flux. Perhaps over time people will come to treat generative AI the same as tools like auto-complete or spellcheck (a possibility we discussed in Section 4). If so, process-based objections may fade away. Of course, the opposite could happen as well: as generative AI becomes more pervasive people may seek out communities of collective refusal. In order to support community self-determination, we should design platforms that allow for both possibilities.
Additionally, generative AI will continue to change, and in time the content-based objections that we heard may no longer be relevant. For example, we consider objections based on AIGC’s accuracy and prose quality, as well as those based on unethical model training processes, to be potentially addressable by technological advances. If online communities are to remain relevant, moderators will need to stay abreast of their members’ stance on this evolving technology; ultimately, if moderator and member perspectives on AIGC differ too much, Reddit’s distributed moderation system allows users to “agree to disagree” by leaving and finding other, more aligned communities (grimmelmann2015virtues). At the time of writing, some platforms have begun incorporating generative AI into their products (lunden_linkedin_2023), likely in an effort to increase engagement. Though this might be aligned with platforms’ profit motives, our interviews suggest it is not aligned with the preferences of all users or communities, as the low effort required to use these features will likely increase the amount of AIGC on the platforms. As we mention in Section 4.1.4, generative AI might bring certain benefits, but given the strong objections that we’ve seen, attempts by platforms to integrate generative AI should be done via a collaborative design process involving users (Gilbert20; Matias16; Matias18). We suggest that platforms that host a variety of community types, such as Reddit, give individual communities the choice to opt-out of generative AI features. If communities choose to opt out, this decision could be communicated to posters via the UI, perhaps by displaying the tools in a disabled state with an informative tooltip. Signaling a community’s stance on AIGC in this way can have the added benefit of additionally discouraging posting AIGC produced off-platform. Additionally, for platforms designed for just one of the community types that we discussed, we recommend considering this community type’s specific values when deciding whether to add generative AI features. Otherwise, platforms run the risk of alienating users (Smith23).
@@ 5.2. The Challenges of Establishing and Enforcing Norms
Our findings relate to and reflect on research about the power of norms to shape behavior in online communities (chancellor18; Chandrasekharan18; kiesler11; Lessig99). We explore the specific challenges of establishing and enforcing community norms about AIGC and discuss ways in which platform designers might be able to aid community moderators with these tasks.
@@ 5.2.1. Clarifying Norms in a Moment of Flux
Our interviews show that self-governing online communities are adapting to AIGC by clarifying norms and codifying these into rules. Prior research has shown that online communities often make rules in response to sudden changes (SeeringEtal2019mec) and that this response can be an effective way to navigate challenging transitions (Kiene16). Moderators sometimes struggle to identify AIGC, which makes enforcing rules about it a challenge; even so, discussing and clarifying rules can help inform the behavior of well-intentioned community members.
Community norms around AIGC are in flux and likely will be for some time. Community members and moderators may have different understandings of whether or not AIGC is acceptable, and this misalignment (Koshy23) can result in well-intentioned users inadvertently violating community norms. Communities can lessen the chance of misalignment by making norms more explicit and known to their members, which has been shown to be effective with other moderation challenges (matias2019preventing). This approach is supported by participants’ reports that AIGC has decreased since they enacted rules banning it, although these reports are based on the moderators’ perceptions, and have not been independently verified. Platform designers can help by updating UIs to ensure that well-meaning posters are aware of community rules. For example, platforms could update their authoring UIs to display a community’s rules, perhaps with a checkbox that users must mark to indicate that they have read and will abide by the rules, similar to how privacy policies are commonly presented. Another option would be to display the rules as placeholder text in the UI element where posts are written, which may remind users of norms at the time of posting (Kiene16). That said, displaying rules in a creative way may not be enough, as past research has demonstrated that privacy policies are rarely read (obar_biggest_2020) and many posters simply do not read a community’s rules before posting (JhaverAppling19). Communities may also consider more thorough “onboarding processes” to ensure new members are familiar with community rules, such as an interview with an existing member. However, this approach adds an additional time burden and might deter potential members who are wary of such a commitment.

### v7

@@ 4.1.3. Governance Concerns
Participants also objected to AIGC because it made the job of moderating their communities more difficult. Eight participants said that they saw malicious uses of AIGC and were concerned that existing moderation techniques would struggle to handle it. These moderators spoke of pre-existing problems, such as spam and harassment, that are more difficult to manage when they are amplified in scale or sophistication by AIGC. While AIGC may be less common than other moderation challenges, it can still be quite disruptive:
@@
Five participants, who moderated the largest subreddits, spoke of AIGC being used in broad attacks on their communities. These attacks often involved automated accounts that post large amounts of spam to disrupt community functioning. One moderator discussed having to ask Reddit staff for help:
@@
Other times these attacks aim to deceive users as a part of influence operations. A moderator of r/news, one of the platform’s largest subreddits, shared a story of one such operation, which they described as, “a highly sophisticated deployment, an attempt to sway a narrative by utilizing accounts that had had their karma bolstered by automated commenting.” This participant references Reddit karma666https://support.reddithelp.com/hc/en-us/articles/204511829-What-is-karma, a reputation points system that communities can use to limit who can post, noting that AIGC can be used to bypass this system through a practice known as karma-farming, in which automated accounts make frequent posts in order to rapidly accumulate karma. Karma-farming is not a new practice, but this moderator thought that AIGC made it harder to detect, and thus more effective. Beyond karma-farming, we also heard of other forms of deception that used AIGC:
@@
Finally, participants expressed concerns about AI making it easier to post objectionable content. One participant who moderated r/Zelda shared that they had seen AI used to produce specific types of objectionable content that their community had already banned, such as commercial or NSFW content. They noted that while this type of content was not new, “it seemed to be new territory for people to expedite these things via AI.”
@@ 4.1.4. Potential Benefits of AIGC
Three moderators mentioned that AI could help non-native English speakers improve their posts. A moderator of r/changemyview reported that, “AI as a tool is going to be useful…it’s useful for people who are poor writers, useful for people who aren’t strong English speakers. And when they need help, AI can fill in those gaps,” though they noted that it was important that such a contribution was still, “sufficiently their words”. Of course, the definition of “sufficient” will be context dependent and may change over time as norms around AI use evolve.
One last benefit shared by our participants was that AIGC could increase engagement in an inactive conversation. For example, one moderator speculated about the potential benefits of AIGC for platforms and provided a hypothetical example:
@@ 4.2. Findings: Online Communities’ Responses to AIGC (RQ2)
To counter the perceived threats of AIGC, moderators have enacted specific changes in their communities. First, our participants emphasized collaborating with their community members to clarify norms about the use of AIGC (Section 4.2.1 below). Once clarified, these norms could be codified into rules, which moderators were then tasked with enforcing (Section 4.2.2).
@@ 4.2.1. Clarifying Community Norms
Almost all participants mentioned eventually codifying community norms into public subreddit rules. Some communities used rules that explicitly banned AIGC or bots, while others used more general rules (such as rules against plagiarism) as effective AIGC bans. Participants shared mixed results about the efficacy of these rules. Seven participants said that the volume of AIGC had decreased in response to new rules, including a moderator of r/AskHistorians who said, “People are using [AIGC] a little bit less now that they know what the rules are. They know that people don’t necessarily want it.” However, a moderator of r/explainlikeimfive expressed skepticism about the efficacy of rules: “I think it may stop some of the average users (to the extent that they actually read the rules, which is a reddit-wide problem.) Bots obviously don’t read rules.” As this participant suggested, explicit rules can guide the behavior of good-faith users, but might not deter bad actors.
@@ 4.2.2. Enforcing Community Rules
Our participants used a variety of strategies to make enforcement of these rules less burdensome. Seven moderators mentioned sharing the work with community members, like a moderator of r/news who noted, “The first line of defense is the users themselves.” However, as mentioned in Section 4.1.2, this runs the risk of spurring false accusations and incivility between members.
Four moderators spoke of their reliance on platform tools, especially automoderator777https://www.reddit.com/wiki/automoderator/, a programmable bot that scans and performs actions on all new posts. Such automated tools lessened the manual effort required by moderators, but they are only as effective as the detection heuristics programmed into them, which, as we discuss in the next section, can be flawed. Moderators said that such tools are not enough to catch all AIGC, and in cases of extreme post volume, such as during coordinated bot attacks, moderators sometimes sought help from Reddit platform admins:
@@ 4.3. Findings: Methods for Detecting AIGC (RQ2)
Our participants shared custom heuristics (Section 4.3.1) and technical solutions (Section 4.3.2) that they use to detect AIGC. Even if participants thought their detection techniques mostly work now, they worry that their tools will be less effective as generative AI improves:
@@ 4.3.1. Detection Heuristics
Our participants reported that these heuristics worked, but also described them as time-consuming. Participants thus turned to technical tools to speed up the detection process.
@@ 4.3.2. Technical Tools
Though they acknowledged such tools’ unreliability, most participants used automated approaches to detect AIGC, including both third-party detection tools and community-developed solutions. Participants spoke of programming the detection heuristics mentioned above into tools such as automoderator, which apply the heuristics to every new post and flag anything suspect. Six participants mentioned manually copying posts into third-party detection tools to see if they were AIGC. A moderator of r/writingprompts reported that they had tested, but never deployed, a homespun bot that would apply third-party detectors automatically to all new posts.
@@ 5. Discussion
Our qualitative investigation offers two main research contributions: (1) a description of moderators’ attitudes towards AIGC in subreddits that restrict its use and (2) insights into how these moderators and their communities have responded to the early challenges of AIGC. On the whole, our findings align with those of recent studies into the impact of generative AI on online communities (Burtch et al., 2024; Wei and Tyson, 2024), which emphasize that the social value of online communities is more important than ever in the age of generative AI. As a company, Reddit seems to agree: in a May 2025 update to the community (u/spez, 2025) Reddit’s CEO shared a similar sentiment:
Reddit works because it’s human. It’s one of the few places online where real people share real opinions. That authenticity is what gives Reddit its value. If we lose trust in that, we lose what makes Reddit…Reddit. Our focus is, and always will be, on keeping Reddit a trusted place for human conversation.
Platform designers and community stewards should both focus on ways to increase this social value, and one way to do this is by empowering communities to take their own stances on AIGC.
As generative AI usage grows, we place our work in dialog with Seering’s call to the CSCW research community to center community self-moderation (Seering, 2020), especially their second “guiding question”: “What are the processes of context-sensitivity in online community moderation, and how might they be better supported?” We believe our findings demonstrate that allowing communities to make context-sensitive decisions about AIGC is central to their self-determination and should be an important topic for CSCW research. Unwanted AIGC in online communities is an active and growing (Lloyd et al., 2025) concern, with Reddit’s CEO describing it as “the worry I hear most often these days from users and mods alike.” (u/spez, 2025) To offer design implications that may help communities address this concern, we expand below on the values that shape community attitudes towards AIGC as well the importance of clear and enforceable community norms.
@@ 5.1. Community Values Shape Attitudes Towards AIGC
In a recent quantitative study (Lloyd et al., 2025), we demonstrated that rules about AI are most common in certain types of subreddits, like those that are larger and those devoted to Content Generation. Our current study digs deeper into the motivations behind these rules and reveals a spectrum of moderator attitudes that vary with the purpose and size of communities. We expect some objections to AIGC will lessen over time, while others will be more persistent.
@@ 5.1.1. Community Types and AIGC Attitudes
Our participants’ concerns about AIGC were largely based on the types of communities they moderated. Accordingly, it is useful to consider our results through the lens of several online community types from the Social Computing literature, such as knowledge communities (Jenkins, 2006), empathic communities (Preece, 1999), communities of practice (Lave and Wenger, 1991), and the five Community Archetypes presented by Prinster et al. (Prinster et al., 2024): Topical Q&A, Learning & Perspective Broadening, Social Support, Content Generation, and Affiliation with an Entity.
We heard much about AIGC’s threat to the social value of a community (Section 4.1.2), which we suspect is most concerning to communities that emphasize authenticity (Smith et al., 2023), like empathic communities (Preece, 1999), or Social Support communities. Interestingly, our prior quantitative study (Lloyd et al., 2025) found that Social Support communities are negatively associated with having a rule about AI. Explicit rules may be less likely in these communities, but our qualitative findings suggest that this should not be interpreted as a lack of concern. Participants who moderated knowledge communities (Jenkins, 2006) like r/AskHistorians and r/explainlikeimfive, which fit the Topical Q&A archetype, also shared their concerns about AIGC’s impact on the social value of their communities. However, these participants additionally discussed concerns about the inaccuracy of AIGC (Section 4.1.1). Arguably the clearest objections came from moderators of communities of practice (Lave and Wenger, 1991) like r/itookapicture and r/WritingPrompts, which fit the Content Generation archetype. These communities shared content quality and social dynamics concerns about AIGC, which have additionally been documented in other recent work (Matatov et al., 2025). While long-standing communities of practice may struggle to adapt to AI, we expect that others will emerge specifically devoted to AI use888https://www.reddit.com/r/aiArt/. Moderators of larger, more impersonal communities that fit the Learning & Perspective Broadening archetype, such as r/news, r/CryptoCurrency, and r/politics, primarily voiced governance concerns (Section 4.1.3). We suspect this is because the size of these communities make them a target of bad actors. Finally, we primarily heard ideological objections from the communities in our sample that fit the Affiliation with an Entity archetype. We suspect this is because in our sample these are Content Generation communities (r/WoW, r/Zelda), and it remains to be seen if communities affiliated with non-media entities, like geographical places or organizations, would share similar concerns.
On the whole, our participants rejected AIGC because they felt the main purpose of their communities was to share posts created by humans—this was true across all of the community types that we encountered (of course, remember that our sampling strategy focused on communities with concerns, not representativeness). This sentiment was made clear by the recent research ethics controversy on r/changemyview, where community members were outraged to learn that researchers had been posting undisclosed AIGC (r/changemyview, 2025).
@@ 5.1.2. Attitudes May Shift Over Time
Our findings indicate that community norms about AIGC are still in flux. Perhaps over time people will come to treat generative AI in the same way they treat tools like auto-complete or spellcheck (a possibility we discussed in Section 4). If so, process-based objections may fade away. Of course, the opposite could happen as well: as generative AI becomes more pervasive people may seek out communities of collective refusal. In order to support community self-determination, we should design platforms that allow for both possibilities.
Additionally, generative AI will continue to change, and in time the content-based objections that our participants shared may no longer be relevant. For example, we consider objections based on AIGC’s accuracy and prose quality, as well as those based on unethical model training processes, to be potentially addressable by technological advances. If online communities are to remain relevant, moderators will need to stay abreast of their members’ stance on this evolving technology; ultimately, if moderator and member perspectives on AIGC differ too much, Reddit’s distributed moderation system allows users to “agree to disagree” by leaving and finding other, more aligned communities (Grimmelmann, 2015). At the time of writing, some platforms have begun incorporating generative AI into their products (Lunden, 2023), likely in an effort to increase engagement. Though this integration might be aligned with platforms’ profit motives, our interviews suggest it is not aligned with the preferences of all users or communities, as the low effort required to use these features will likely increase the prevalence of AIGC. As we mention in Section 4.1.4, generative AI might bring certain benefits, but given the strong objections that we have seen, attempts by platforms to integrate generative AI should be done via a collaborative design process involving users (Matias and Mou, 2018). We suggest that platforms that host a variety of community types, such as Reddit, give individual communities the choice to opt-out of generative AI features. If communities choose to opt-out, this decision could be communicated to posters via the UI, perhaps by displaying the tools in a disabled state with an informative tooltip. Signaling a community’s stance on AIGC in this way can have the added benefit of discouraging the use of AIGC that is produced off-platform. Additionally, for platforms designed for just one of the community types that we discussed, we recommend considering the community type’s values when deciding whether to add generative AI features. Otherwise, platforms run the risk of alienating users (Smith et al., 2023).
@@ 5.2. The Challenges of Establishing and Enforcing Norms
Our findings relate to and reflect on research about the power of norms to shape behavior in online communities (Chancellor et al., 2018; Chandrasekharan et al., 2018; Kiesler et al., 2011; Lessig, 1999). We explore the specific challenges of establishing and enforcing community norms about AIGC and discuss ways in which platform designers might be able to aid community moderators with these tasks.
@@ 5.2.1. Clarifying Norms in a Moment of Flux
Our interviews show that self-governing online communities are adapting to AIGC by clarifying norms and codifying these into rules. Prior research has shown that online communities often make rules in response to sudden changes (Seering et al., 2019) and that this rule-making can be an effective technique for navigating transitions (Kiene et al., 2016). Moderators sometimes struggle to identify AIGC, which makes enforcing rules about it a challenge; even so, discussing and clarifying rules can help inform the behavior of well-intentioned community members.
As noted above, community norms around AIGC are in flux and likely will be for some time. Community members and moderators may have different understandings of whether or not AIGC is acceptable, and this misalignment (Koshy et al., 2023) can result in well-intentioned users inadvertently violating community norms. Communities can reduce misalignment by making their norms more explicit and known to members, which has proven effective with other moderation challenges (Matias, 2019b). Indeed, our participants’ report that AIGC has decreased since they enacted bans, though these trends have not been independently verified and may change over time. Platform designers can help by updating UIs to ensure that well-meaning posters are aware of community rules. For example, platforms could update their authoring UIs to display a community’s rules, perhaps with a checkbox that users must mark to indicate that they have read and will abide by the rules, similar to how privacy policies are commonly presented. Another option would be to display the rules as placeholder text in the UI element where posts are written, which may remind users of norms at the time of posting (Kiene et al., 2016). That said, displaying rules in a creative way may not be enough, as past research has demonstrated that privacy policies are rarely read (Obar and Oeldorf-Hirsch, 2020) and many posters simply do not read a community’s rules before posting (Jhaver et al., 2019a). Communities may also consider more thorough “onboarding processes” to ensure new members are familiar with community rules, such as an interview with an existing member. However, this approach adds an additional time burden and might deter potential members who are wary of such a commitment.